OpenStudy (annie96):

2. 5 ( 3 – m) + 5m = 15 3. -2 |x| - 7 = -3

OpenStudy (annie96):

1.) 4-2y=2(9-y) 2y=2+(-18y) 2y=-16y y=-8

OpenStudy (annie96):

no

OpenStudy (zpupster):

2(9-y )= 18-2y by distributive property

OpenStudy (bibby):

Expand it out using the distributive property as zpup just did 5( 3 – m) + 5m = 15 (5*3)-(5*m)+(5m) = 15

OpenStudy (anonymous):

The three equations are trick equations. 4 - 2 y = 2 (9 - y) 4 - 2 y = 18 - 2 y 4 != 18 5 (3 - m) + 5 m == 15 15 - 5 m + 5 m == 15 15 = 15 -7 - 2 Abs[x] == -3 Abs[x] != -2

OpenStudy (anonymous):

Sorry. == is = in Mathematica. Did not edit the last two solutions.

OpenStudy (bibby):

Just so we're clear != == =/= || ≠:)

OpenStudy (anonymous):

!= "not equal" == "equal"

OpenStudy (bibby):

this was programming humor yes thank

OpenStudy (anonymous):

@bibby Ha Ha. Very clever. I happen to be a former programmer.

OpenStudy (annie96):

How do you solve number 3.

OpenStudy (anonymous):

-7 - 2 Abs[x] = -3 - 2 Abs[x] = 7-3 - 2 Abs[x] = 4 Abs[x] = 4/(- 2) Abs[x] ≠ -2